# A Non-perturbative General Approximation Scheme (NGAS) for Interacting Quantum Systems with Perturbation Theory for Arbitrary Strength of Interaction

### v2

@@ VII.3 The Quartic Double Well Oscillator (QDWO)
The QDWO is also an extensively studied system [94] because of its theoretical importance and practical applications. The Hamiltonian of the system is given by:
@@
The crucial  -ve   sign of the $x^{2}$- term generates a quite different physical situation than the case of the QAHO, even in the classical limit. The ‘classical’ potential, ${}V_{c}\equiv-\frac{1}{2}x^{2}~{}+~{}gx^{4}~{}$ exhibits the familiar double-well shape with symmetric minima. These minima are located at positions ±${}\pm\frac{1}{2\sqrt{g}}~{}$ and each with depth ${}\frac{1}{16g}~{}$. As ${}g~{}$ becomes smaller and smaller the depth of the two- wells become deeper and deeper. The actual low lying energy eigen-states of the problem become radically different from the trial wave function offered by the harmonic basis. This fact severely handicaps the convergence of the resultant perturbation theory. In principle, the natural solution would be the simultaneous use of two harmonic basis-centers around the positions of the minima at ±${}\pm\frac{1}{2\sqrt{g}}~{}$. However, the implementation of that idea, although possible, implies the use of nonorthogonal states which is rather cumbersome. The   other  difficulty is that the theory is not defined for ${}g\rightarrow 0~{}$, because the ground state does not exist in that limit due to the non-existence of a lower limit to $V_{c}$. In that sense,  the SHO is not the free-field limit of the QDWO. Therefore, the ’standard formulation of perturbation theory (SFPT)’ is not applicable as such, to this case. The perturbation expansion of the eigenvalues ${}E_{n}(g)~{}$ in powers  of ${}g~{}$ is divergent [95-99] for all ${}g~{}>~{}0~{}$. This fact may be understood qualitatively by noting that the addition of the term ${}gx^{4}~{}$ turns a completely continuous eigenvalue spectrum of ${}p^{2}~{}-~{}x^{2}~{}$ into a completely discrete spectrum bounded from below. A non perturbative treatment is therefore, necessary. The WKB method is one such method which is well suited especially for lower energy eigenvalues.
Moreover, it can be mentioned that the case of the QDWO is not Borel-summable [95-99] in SFPT for any value of the coupling strength. Several modern developments such as the theory of resurgence and trans-series [95-97],distributional-Borel summation [98], generalized Borel-Pade method [99] etc. are basically motivated to surmount this problem. In view of the above scenario, the case of the QDWO assumes special relevance for investigation in NGAS.
However, as shown below, in contrast to the situation in SFPT the NGAS can be successfully applied to the case of the QDWO, even in the LO, using completely analogous procedure as in the case of the QAHO, for a considerable larger range of values of ${}n~{}$ and ${}g~{}$.
To develop the NGAS for QDWO the “ approximating Hamiltonian (AH)”for the system is again chosen in the harmonic approximation as:
@@
but generalized to take into account, the spontaneous symmetry breaking (SSB), through a nonzero vacuum expectation value for $x$ denoted as $\sigma$. Accordingly, the various average values, analogous to eqs.(27a-c) are now given by
@@
Eqs.(38-40) enable the evaluation of $\langle H\rangle$ in terms of the input parameters. These parameters: $\omega$, $\sigma$ and $h_{0}$ are then determined analogously as in the case of the AHO, in terms of $g$ and $\xi$ However, a distinct feature in the case of the QDWO is the occurrence of “quantum-phase transition(QPT)” [82] governed by a “critical coupling” $g_{c}(\xi)$ given by the expression:
@@
such that the “Spontaneously Symmetry Broken)(SSB)” phase is realized with $\sigma\neq 0$ for $g\leq g_{c}(\xi)$, whereas the “symmetry-restored (SR) phase” is obtained with $\sigma=0$ when $g>g_{c}(\xi)$. (Numerically, we have: $g_{c}=0.0907218$ , for the ground state of the DWO).
The transition across $g=g_{c}(\xi)$ being discontinuous, the two phases are governed by distinct expressions for $\omega$ and $E_{0}$, which are not analytically connected. It is, therefore, necessary to consider the two phases separately. However, owing to the rather small value of $g_{c}$ the “SSB phase” exists only over a very limited range of $g$ ,i.e when $0\leq g_{c}\leq 0.0907218$ for the ground state and $g_{c}$ takes still smaller values for the higher excited states, eventually vanishing for large $n$ We therefore, confine here to reporting the results for the “SR phase” only.
In this phase, we have $g>g_{c}(\xi)$; $\sigma=0$; $\omega$ satisfies the gap equation
@@
The energy levels in this phase, are given by the simple expression:
@@
and $h_{0}=E_{0}-\omega\xi$. In eqn.[42] $\omega$ is the solution of eqn.[41] which can be obtained in analogous manner and is given by
@@
where, ${}\rho^{-1}~{}=~{}243g^{2}f^{2}(\xi)~{}$. As in the case of the QAHO, the LO-results $E_{0}$ captures the dominant contribution in reproducing the energy spectrum [8,82] to within a few percent (For the ground state energy, see, TABLE-I and -II in Section-X). In these Tables, we also display the improvement of the LO-results obtained by higher order corrections in the MFPT which is discussed in Section-X .
@@ VII.4 The Sextic Anharmonic Oscillator (SAHO)
The sextic AHO-system is an example of higher anharmonicity, which is also widely investigated [100]. This system is interesting and important in its own right as it finds application in diverse areas of physics [100]. As a theoretical laboratory, this system together with its double-well counterpart provide perhaps the simplest examples, for which supersymmetric quantum mechanics (SUSYQM) has definite predictions [101] for the energy-levels for specific values of the Hamiltonian-parameters. Hence, various models and approximation methods can be tested against the exact analytic results of SUSYQM. Besides, owing to the higher anharmonicity, the divergence of the SFPT at higher-orders becomes even more severe [102], $E_{n}\sim\Gamma[n(m-1)]$ for $gx^{2m}$ type of AHO. This result endows the sextic-AHO with added importance for testing convergent-approximation methods. It may also be noted that the“Wick-ordering” method of Caswell [27] is not directly applicable to this case since the method generates $x^{4}$ counter-terms which are not present in the original (non-ordered) Hamiltonian. For the above stated reasons of practical applicability and theoretical importance, the sextic-AHO provides a unique testing ground for the NGAS, as described in the following.
The Hamiltonian for the system is given by
@@
Results for the SAHO can be obtained by following analogous procedure as in the case of QAHO along with the following input [82] for $\omega$, $h_{0}$ and $E_{0}$: $\omega$ is given by the real, positive root of the ’gap-equation’
@@
One then obtains [82] the following simple expression for the energy-levels of the sextic AHO in the LO:
@@
Again, the LO-results for the energy are obtained accurately [82] to within a few percent over a wide range in $\xi$ and for arbitrary physical values of $g\geq 0$. Further improvement in accuracy is achieved by application of the MFPT as discussed in Section-X .
@@ VII.5 The Octic- Anharmonic Oscillator (OAHO)
The octic-anharmonic oscillator, like its quartic- and sextic-counterparts, finds applications in modelling molecular physics,lattice-vibrations in solids and in quantum chemistry. Because of the higher anharmonicity, the system also provides the theoretical laboratory for more stringent-tests for non-perturbative approximation schemes in quantum theory since the divergence of the SFPT at higher orders of perturbation, becomes still more severe [102] in this case.
The system is considered here to test the generality, uniform applicability and the reliability of NGAS to the case of higher anharmonicities in the AHI. For this case, the detailed description follows:
The Hamiltonian is given by
@@
We employ in this case, the “approximating Hamiltonian (AH)” given by,
@@
Eqn.[46] corresponds to “shifted” effective Harmonic oscillator. The energy spectrum , as in the earlier cases, is
@@
where, as before:    $\xi~{}=~{}(n+1/2)~{}$ ;${}n~{}=~{}0,1,2,.....$ It is next required to determine the frequency ‘$\omega$’ and ‘$h_{0}$’. For this purpose, we note that the quantum average of the eqn.(45) is given by:
$<H>~{}=~{}<H_{0}>~{}=~{}\frac{1}{2}<p^{2}>+\frac{1}{2}<x^{2}>+g<x^{8}>$ (48)
where the quantum-averages on the r.h.s. of eqn.(48) can be calculated using the standard properties of creation-/annihilation operators as in earlier cases (see, eqn.(27)). After applying the variational minimisation condition and considering $\sigma=0$ to get “physically acceptable” solution, we are led [82] to the simplified GE, given by:
@@
The solution of eqn.(49) determines the frequency ‘$\omega$’ of the “shifted” harmonic oscillator. To obtain the energy levels one substitutes ${}\sigma~{}=~{}0$ and ‘$\omega$’ as the solution of eqn.(49). This leads, after some simplification, to the following simple formula[82]:
@@
where ‘$\omega$’ is obtained by solving eqn.(49) numerically. We compare the LO- result in NGAS [82] with earlier computations [103] over a wide range of values of ‘$g$’ and `$n^{\prime}$. It can be seen from this comparison [82] that the results obtained in the LO of NGAS are already quite accurate over the full range of the parameters, which demonstrates the generality of the method and uniformity of the approximation with increasing anharmonicity.
As stated earlier, an important aspect of NGAS is its flexibility for accommodating different choices for the input Hamiltonian, AH in order to compare and improve up on the final results for the system-Hamiltonian. We demonstrate this feature in the next Section.
@@ VIII Flexibility of Input in NGAS - The square-well approximation
In the previous Sections, the ’simple harmonic-approximation’ was considered as the natural choice for the input/ approximating Hamiltonian (AH) in NGAS for dealing with different cases of AHI. In this Section, we exploit the freedom of choice of the AH in NGAS to consider the same corresponding to the infinite-square-well (ISW)-potential. One important motivation in selecting the ISW-potential as the input, is the pedagogical aspect, as well as the simplicity of this system. It is well-known that the ISW-potential constitutes one of the simplest systems admitting exact analytical solution. As such, it is included in any standard course of introductory quantum mechanics. The establishment of an approximation connecting the AHO/DWO system to the ISW-case, may therefore be considered as an illustrative example of application of a standard text-book topic to advanced research.
Admittedly, however, the ISW- approximation is perhaps the crudest among possible choices of AH. This is due to the fact that the system described by the ISW potential propagates freely between the infinite walls, in stark contrast to the actual situation for the AHO/DWO. Nevertheless, we choose it on purpose here, in order to test the robustness and tolerance of NGAS to the crudest possible input approximation, which could be yet another motivation.
Moreover, as a by-product of this study, it is possible to obtain an approximation for the celebrated case of the SHO in the limit as the anharmonic- coupling strength is set to zero in the AHO-Hamiltonian, thus gaining further insight into the accuracy of NGAS.
In the next subsection, we demonstrate the method by applying it to the case of the quartic-AHO choosing the ISW-Hamiltonian as the input. The LO-results for the energy spectrum are obtained and compared with the results from other calculations. We also highlight the non-perturbative aspects and the analytic structure of energy as a function of the quartic-coupling.
@@ VIII.1 The Quartic Anharmonic Oscillator in ISW-approximation
The quartic-AHO is defined by the Hamiltonian earlier (see, eqn.(24)) and reproduced below (with a different equation-number for ease of reference):
@@
where, $g>0$
The input-Hamiltonian for the ISW-potential is defined by:
@@
where the potential, $V_{SW}$ is given by:
${V_{SW}(x)}~{}=~{}\left\{\begin{array}[]{rc}\infty&\mbox{$|x|\geq a$}\\ h&\mbox{$~{}~{}|x|<~{}a~{}$},\\ \end{array}\right.$ (55)
@@
is easily solved subject to the boundary-condition:
$\phi^{SW}_{n}(x)~{}~{}=~{}0~{}~{}~{}~{}~{}\textrm{for}~{}~{}|x|\geq{a},$ (57)
which ensures the physical requirement of absolute confinement of the system between the (infinite) potential barriers. For ${}~{}|x|<a,$ the normalized eigenfunctions, vanishing on the boundary of the potential-well, are given by:
$\phi^{SW}_{n}(x)\equiv\phi^{(-)}_{n}(x)~{}=~{}\frac{1}{\sqrt{a}}~{}sin\left(\frac{n\pi x}{2a}\right);~{}~{}\textrm{for}~{}~{}$
@@
In the above equations, the (±)plus-or-minus($\pm$)( ± ) super-scripts correspond to the even(odd)-parity solutions. The energy eigenvalues are trivially obtained by solving the Schrödinger equation and given by
@@
[…]
The evaluation of $<H>$ using the above definition,eqn.[59]is straight-forward. Noting that:
@@
each term in eqn.(60) can be computed easily by exploiting parity-invariance, which forbids parity-changing transitions, i.e. ${}<\phi_{n}^{+}|H|\phi_{n}^{-}>=0=<\phi_{n}^{-}|H|\phi_{n}^{+}>$.
@@
The minimization of the expression for $<H>$ as given above, with respect to $u\equiv~{}\left(1/a^{2}\right)$ leads to the following equation:
@@
where, $u$ is given by eqn.(66). The remaining parameter, `$h^{\prime}$ can then be determined by substitution of eqn.(68) in eqn.(58), given by:
@@
At this point, several observations are in order:
(i) note that the free parameters of the input-Hamiltonian,i.e. `$a^{\prime}$ and `$h^{\prime}$ acquire functional dependence on `$g^{\prime}$ and `$n^{\prime}$ through eqn.(66) and eqn.(69). This in turn , implies that the input-Hamiltonian,$H_{0}^{SW}$ (see,eqn.(52) also becomes a function of $g$ and $n$. An obvious consequence is that the eigenfunctions of $H_{0}^{SW}{(g,n)}$ corresponding to different eigenvalues become non-orthogonal as the latter now depends on the excitation-label ’n’,i.e.
@@
(ii) It is seen from eqn.(66) and eqn.(68) that the energy in the LO is a non-analytic function of $g$ at the origin. Hence, the results expressed in these equations are not accessible to ordinary perturbation theory as a power-series expansion in $g$. In this sense, the LO-results are non-perturbative.
(iii) The cube-root singularity at $g=0$ and the branch-point-structure in the $complex~{}g-plane$ as given in eqn.(68) are in conformity with rigorous derivation [11] of the analytic structure of the energy of the quartic-AHO in the coupling strength-plane using sophisticated tools of complex-analysis. It is remarkable that, the correct analytic-structure in $g$ arises here as a simple consequence of the CEQA and the CO of NGAS( see, eqn.(16) and eqn.(17)). (iv) Further, it needs to be emphasized here that the observed features noted above under (i)-(iii) also hold in the analogous case of the simple-harmonic approximation [82] applied to the same system of the QAHO/QDWO as described here in previous Sections. This fact demonstrates that these features are the inherent properties of the scheme, NGAS, being independent of the choice of the input Hamiltonian.
[…]
It is seen from this tabulation that the LO-results are uniformly accurate to within $\sim(2-12)~{}\%$ of the standard results over the full range of $n_{s}$ and $g$. It is further seen that the accuracy of the LO-approximation increases significantly by the inclusion of the 2nd-order correction in IPT.
@@
$n_{s}$ $g$ $E_{n}^{LO}$ $Error(LO)$ $E_{n}^{(2)}$ $Exact$ $Error$
[…]
@@
The energy-eigenfunctions in the LO are those as given by eqn.(56) and eqn.(57) but with the width-parameter,`$a^{\prime}$ appearing in these equations now acquire functional dependence on $g$ and $n$ in accordance with eqn.(66).
We discuss the LO-approximation for the SHO in the next subsection.

### v3

@@ VII.3 The Quartic Double Well Oscillator (QDWO)
The quantum double-well oscillator (QDWO) is a system of significant theoretical and practical interest [94]. The Hamiltonian of this system is expressed as:
@@
The negative sign of the $x^{2}$ term introduces a distinct physical scenario compared to the quantum anharmonic oscillator (QAHO), even in the classical limit. The classical potential, $V_{c}\equiv-\frac{1}{2}x^{2}~+~gx^{4}~$ displays the characteristic double-well shape with symmetric minima. These minima are located at $\pm\frac{1}{2\sqrt{g}}~$, each with a depth of $\frac{1}{16g}~$. As $g~$ decreases, the depth of the wells increases. The actual low-lying energy eigenstates of the system differ significantly from those predicted by the harmonic basis, which impedes the convergence of the resulting perturbation theory. Ideally, a natural solution would involve the simultaneous use of two harmonic basis centers around the minima at $\pm\frac{1}{2\sqrt{g}}~$. However, implementing this approach requires the use of nonorthogonal states, which is cumbersome. Another challenge is that the theory is not defined for $g\rightarrow 0~$, as the ground state does not exist in this limit due to the absence of a lower bound for $V_{c}$. Consequently, the simple harmonic oscillator (SHO) is not the free-field limit of the QDWO. Therefore, the standard formulation of perturbation theory (SFPT) is not applicable in this case. The perturbation expansion of the eigenvalues $E_{n}(g)~$ in powers of $g~$ is divergent [95-99] for all g>0~g~>~0~. This divergence can be qualitatively understood by noting that the addition of the $gx^{4}~$ term transforms a completely continuous eigenvalue spectrum of $p^{2}~-~x^{2}~$ into a discrete spectrum bounded from below. A non-perturbative approach is thus necessary. Furthermore, it is noteworthy that the QDWO case is not Borel-summable [95-99] in SFPT for any value of the coupling strength. Recent advancements, such as the theory of resurgence and trans-series [95-97], distributional-Borel summation [98], and the generalized Borel-Padé method [99], have been developed to address this issue. In light of these considerations, the QDWO case assumes special relevance for investigation in the  non-perturbative general approximation scheme (NGAS).
In contrast to the situation in SFPT, the NGAS can be effectively applied to the case of the QDWO, even in the LO, using a procedure analogous to that employed for the QAHO, over a considerably larger range of values for $n~$ and $g~$. To develop the NGAS for QDWO, the ”approximating Hamiltonian (AH)” for the system is again selected within the harmonic approximation as follows:
@@
but generalized to account for spontaneous symmetry breaking (SSB) through a nonzero vacuum expectation value for $x$, denoted as $\sigma$. Consequently, the various average values, analogous to eqs.(27a-c), are now expressed as:
@@
Equations (38-40) facilitate the evaluation of $\langle H\rangle$ in terms of the input parameters. These parameters: $\omega$, $\sigma$, and $h_{0}$ are then determined analogously to the case of the AHO, in terms of $g$ and $\xi$. However, a distinctive feature in the case of the QDWO is the occurrence of a ”quantum-phase transition (QPT)” [82] governed by a ”critical coupling” $g_{c}(\xi)$, given by the expression:
@@
such that the ”Spontaneously Symmetry Broken (SSB)” phase is realized with $\sigma\neq 0$ for $g\leq g_{c}(\xi)$, whereas the ”symmetry-restored (SR) phase” is obtained with $\sigma=0$ when $g>g_{c}(\xi)$. Numerically, we have: $g_{c}=0.0907218$, for the ground state of the DWO. The transition across $g=g_{c}(\xi)$ is discontinuous, with the two phases governed by distinct expressions for $\omega$ and $E_{0}$, which are not analytically connected. It is, therefore, necessary to consider the two phases separately. However, due to the relatively small value of $g_{c}$, the ”SSB phase” exists only over a very limited range of $g$, i.e., when $0\leq g_{c}\leq 0.0907218$ for the ground state, and $g_{c}$ takes even smaller values for the higher excited states, eventually vanishing for large $n$. Therefore, we confine our discussion to reporting the results for the ”SR phase” only. In this phase, the condition $g>g_{c}(\xi)$ is satisfied, with $\sigma=0$, and $\omega$ fulfilling the gap equation:
@@
The energy levels in this phase are represented by the expression:
@@
and $h_{0}=E_{0}-\omega\xi$. In equation [42], $\omega$ is the solution of equation [41], which can be derived analogously and is expressed as:
@@
where $\rho^{-1}~=~243g^{2}f^{2}(\xi)~$.
Similar to the case of the Quantum Anharmonic Oscillator (QAHO), the leading-order (LO) results for $E_{0}$ effectively capture the primary contribution in reproducing the energy spectrum [8,82] with a precision of a few percent - for the ground state energy, refer to TABLE-I and -II in Section-X.  These tables also illustrate the enhancement of accuracy of the LO results achieved through higher-order corrections in the ‘Mean Field Perturbation Theory (MFPT)’, as discussed in Section-X.
@@ VII.4 The Sextic Anharmonic Oscillator (SAHO)
The sextic anharmonic oscillator system is a prime example of higher anharmonicity and has been the subject of extensive research [100]. This system holds considerable interest and significance due to its applications in various branches of physics [100]. As a theoretical construct, it, along with its double-well variant, represents one of the simplest cases where supersymmetric quantum mechanics (SUSYQM) offers precise predictions [101] for energy levels based on specific Hamiltonian parameters. This allows for the evaluation of various models and approximation techniques against the exact analytical outcomes provided by SUSYQM. Additionally, the increased anharmonicity results in a more severe divergence of the SFPT at higher orders [102], characterized by $E_{n}\sim\Gamma[n(m-1)]$ for $gx^{2m}$ types of AHO. This feature underscores the sextic-AHO’s importance in assessing convergent-approximation methods.  For these practical and theoretical reasons, the sextic-AHO serves as a distinctive testing platform for the NGAS, as detailed below.
The Hamiltonian for this system is defined as:
@@
The results for the SAHO can be obtained by employing a method similar to that used for the QAHO, utilizing the following input [82] for $\omega$, $h_{0}$, and $E_{0}$:  The value of $\omega$ is determined by the real, positive root of the ’gap-equation’,
@@
This leads to the derivation [82] of the following expression for the energy levels of the sextic AHO in the LO:
@@
The LO results for the energy are obtained with high accuracy [82], within a few percent, across a broad range of $\xi$ and for any physical values of $g\geq 0$. The LO results along with perturbative improvement of accuracy through MFPT are discussed in Section-X.
@@ VII.5 The Octic- Anharmonic Oscillator (OAHO)
The octic-anharmonic oscillator, akin to its quartic and sextic analogs, is utilized in the modeling of molecular physics, lattice vibrations in solids, and quantum chemistry. Owing to its enhanced anharmonicity, this system also functions as a theoretical - laboratory for more stringent evaluations of non-perturbative approximation schemes in quantum theory, as the divergence of the SFPT at higher perturbation orders becomes increasingly pronounced [102] in this case. The system is examined here to assess the generality, uniform applicability, and reliability of NGAS in the context of higher anharmonicities in the AHI. For this purpose, the detailed description is as follows:
The Hamiltonian is expressed as
@@
In this context, we utilize the ”approximating Hamiltonian (AH)” defined as follows:
@@
Equation [46] represents a ”shifted” effective harmonic oscillator. The energy spectrum, consistent with previous instances, is given by
@@
where, as previously defined:    $\xi~=~(n+1/2)$~ ;n=0,1,2,..~n~=~0,1,2,......
It is subsequently essential to determine the frequency $\omega$’ and $h_{0}$’. To this end, we observe that the quantum average of equation (45) is expressed as:
$H>~=~<H_{0}>~=~\frac{1}{2}<p^{2}>+\frac{1}{2}<x^{2}>+g<x^{8}>$ (48)
where the quantum averages on the right-hand side of equation (48) can be computed using the standard properties of creation/annihilation operators as in prior cases (refer to equation (27)). By applying the variational minimization condition and considering $\sigma=0$ to achieve a ”physically acceptable” solution, we derive [82] the simplified gap equation, expressed as:
@@
The solution of equation (49) determines the frequency $\omega$’ of the ”shifted” harmonic oscillator. To obtain the energy levels, one substitutes $\sigma~=~0$ and $\omega$’ as the solution of equation (49). This leads, after some simplification, to the following simple formula [82]:
@@
where $\omega$’ is obtained by solving equation (49) numerically. We compare the leading-order result in NGAS [82] with previous computations [103] over a wide range of values of $g$ and $n$. It is evident from this comparison [82] that the results obtained in the leading order of NGAS are already quite accurate over the full range of the parameters. This demonstrates the generality of the method and uniformity of the approximation with increasing anharmonicity.
As previously mentioned, a significant feature of NGAS is its flexibility in accommodating various choices for the input Hamiltonian, AH, to facilitate comparison and enhancement of the final results for the system-Hamiltonian. This feature will be demonstrated in the subsequent Section.
@@ VIII Flexibility of Input in NGAS - The square-well approximation
In the preceding sections, the ’simple harmonic-approximation’ was identified as the natural choice for the input/approximating Hamiltonian (AH) in NGAS when addressing various cases of AHI. In this Section, we explore the flexibility in selecting the AH in NGAS by considering the infinite-square-well (ISW) potential. A significant motivation for choosing the ISW potential as the input is its pedagogical value and the simplicity of this system. It is well-established that the ISW potential represents one of the simplest systems that allows for an exact analytical solution. Consequently, it is included in any standard introductory quantum mechanics course. Establishing an approximation that connects the AHO/DWO system to the ISW case may thus serve as an illustrative example of applying a standard textbook topic to advanced research. However, it must be acknowledged that the ISW approximation is perhaps the crudest among the possible choices of AH. This is because the system described by the ISW potential propagates freely between the infinite walls, in stark contrast to the actual situation for the AHO/DWO. Nevertheless, we deliberately choose it here to test the robustness and tolerance of NGAS to the crudest possible input approximation, which could serve as an additional motivation. Furthermore, as a by-product of this study, it is possible to derive an approximation for the well-known case of the SHO in the limit where the anharmonic coupling strength is set to zero in the AHO Hamiltonian, thereby gaining further insight into the accuracy of NGAS. In the subsequent subsection, we demonstrate the method by applying it to the case of the quartic AHO, selecting the ISW Hamiltonian as the input. The LO results for the energy spectrum are obtained and compared with results from other calculations. We also emphasize the non-perturbative aspects and the analytic structure of energy as a function of the quartic coupling.
@@ VIII.1 The Quartic Anharmonic Oscillator in ISW-approximation
The quartic anharmonic oscillator (quartic-AHO) is characterized by the Hamiltonian previously introduced (refer to equation (24)) and is reiterated here for convenience with a different equation number:
@@
where $g>0$.
The Hamiltonian for the infinite square well (ISW) potential is defined as follows:
@@
where the potential $V_{SW}$ is specified by:
@@
The ISW potential is notably characterized by two free parameters: the width, denoted as $2a$, and the depth, denoted as $h$. Additionally, the potential is selected to be symmetric under spatial inversion, such that $V_{SW}(-x)=V_{SW}(x)$, to ensure consistency with the symmetry of the original Hamiltonian as specified in equation (51). The eigenvalue equation for $H_{0}^{SW}$ is presented as follows:
@@
This equation is readily solvable under the boundary condition:
$\phi^{SW}_{n}(x)~~=~ ~~~~~\textrm{for}~~|x|\geq{a},$ (57)
which ensures the physical requirement of absolute confinement of the system within the infinite potential barriers. For $|x|<a$, the normalized eigenfunctions, which vanish at the boundary of the potential well, are given by:
$\phi^{SW}_{n}(x)\equiv\phi^{(-)}_{n}(x)~=~\frac{1}{\sqrt{a}}~\sin\left(\frac{n\pi x}{2a}\right);~~\textrm{for}$~~
@@
In the above equations, the $(\pm)$ superscripts correspond to the even(odd)-parity solutions. The energy eigenvalues are straightforwardly obtained by solving the Schrödinger equation and are given by
@@
The subsequent task involves determining the adjustable parameters, $\it{a^{\prime}}$ and $\it{h^{\prime}}$. The width parameter $a^{\prime}$ can be determined by the variational minimization of <H>{<H>}. Here, we use the notation: $\hat{A}$>} to define the quantum average/expectation value of the operator, $^{\hat{A}}$ as given below:
$\langle{\hat{A}}\rangle~\equiv~\int_{-a}^{+a}dx~{\phi_{n}^{\ast}(x)~\hat{A}~\phi_{n}(x)}$ (61)
The evaluation of <H><H> using the aforementioned definition, eqn.[59], is straightforward. It is noted that:
@@
each term in eqn.(60) can be computed easily by exploiting parity-invariance, which forbids parity-changing transitions, i.e. $\phi_{n}^{+}|H|\phi_{n}^{-}>=0=<\phi_{n}^{-}|H|\phi_{n}^{+}>$.
@@
The minimization of the expression for <H><H> as given above, with respect to $u\equiv~\left(1/a^{2}\right)$ leads to the following equation:
@@
where, $u$ is given by eqn.(66). The remaining parameter, $h^{\prime}$ can then be determined by substitution of eqn.(68) in eqn.(58), given by:
@@
At this juncture, several observations are warranted:
@@
It follows from eqn. (66) and eqn. (68) that the energy in the LO is a non-analytic function of $g$ at the origin. Consequently, the results expressed in these equations are not amenable to ordinary perturbation theory as a power-series expansion in $g$. In this context, the LO results are non-perturbative.
(iii) The cube-root singularity at $g=0$ and the branch-point structure in the complexgplanecomplex~g-plane, as presented in eqn. (68), align with the rigorous derivation [11] of the analytic structure of the energy of the quartic-AHO in the coupling strength-plane, utilizing advanced tools of complex analysis. It is noteworthy that the correct analytic structure in $g$ emerges here as a straightforward consequence of the CEQA and the CO of NGAS (see, eqn. (16) and eqn. (17)). (iv) Furthermore, it is crucial to emphasize that the observed features noted above under (i)-(iii) also apply in the analogous case of the simple harmonic approximation [82] applied to the same system of the QAHO/QDWO as described here in previous Sections. This fact demonstrates that these features are the inherent properties of the scheme, NGAS, being independent of the choice of the input Hamiltonian.
In Table-I, we present the leading-order (LO) results for the energy levels of the quartic anharmonic oscillator (AHO) at selected values of the oscillator-level index $n_{s}$ and the quartic coupling $g$. It is important to note that the infinite square well (ISW) level index and the oscillator-level index differ by one unit, i.e., $n_{s}=n-1$, for $n=1,2,3,4,\ldots$. Additionally, this table includes the energy levels $E_{n}^{(2)}$, which incorporate second-order corrections obtained through the ’improved perturbation theory’ (IPT) (refer to Subsection-D below), as well as previous results from reference [105], denoted as $Exact$ in column 5, for comparison. The entries in the last column and the third column represent the relative percentage error with respect to the $Exact$ results.
The tabulation reveals that the LO results are consistently accurate within approximately $\sim(2-12)~\%$ of the standard results across the entire range of $n_{s}$ and $g$. Furthermore, it is evident that the accuracy of the LO approximation is significantly enhanced by the inclusion of the second-order correction in IPT.
@@
$n_{s}$ $g$ $E_{n}^{LO}$ $Error(LO)$ $E_{n}^{(2)}$ ExactExact ErrorError
$(\%)$} ref.(8) $(\%)$}
@@
The energy eigenfunctions in the LO are as specified by equations (56) and (57), with the width parameter $a$ in these equations now exhibiting functional dependence on $g$ and $n$ in accordance with equation (66).
The LO approximation for the simple harmonic oscillator (SHO) will be discussed in the subsequent subsection.